A solution A was prepared by dissolving 60 g of NaOH in 100ml of distilled water. Later, 10 ml where taken and poured into a 250 ml volumetric flask. A solution B is obtained

a-calculate mass concentration of solution A


b- Deduce molar concentration of A


c- Find molar concentrantion of B Given : M NaOH= 40g/mol


1
Expert's answer
2021-03-24T01:09:32-0400

a. mass concentration of solution A = 60g*100%/160g = 37.5%

b. n(NaOH) = 60g/40g/mol = 1.5 mol

molar concentration of A = 1.5mol/0.16L = 9.375 M

c. n(NaOH) = 9.375M*0.01L = 0.09375 mol

molar concentrantion of B = 0.09375 mol/0.25L = 0.375 M